OpenStudy (johnnewman):

Robby invested $18,400 at 7% interest. He earned $9016 interest at the end of x years. What is the value of x? A. 10 B. 9 C. 8 D. 7

OpenStudy (jonathan34):

well hello

OpenStudy (johnnewman):

hey how can I work it out

OpenStudy (jonathan34):

well am working on it write know i all give you the answer just wait a sec

OpenStudy (jonathan34):

D

OpenStudy (jonathan34):

X=9016/18400*0.07 X=9016/1288

OpenStudy (jonathan34):

thank you for the medal thing !

OpenStudy (jonathan34):

thank you for 2 other person

OpenStudy (johnnewman):

your welcome

OpenStudy (jonathan34):

ok if you need more help just tell me and thx one more time